# OrionNav: Online Planning for Robot Autonomy with Context-Aware LLM and Open-Vocabulary Semantic Scene Graphs

### v1

@@ V-C1 Frontier-Search
Out of the 9 trials, Frontier-Search failed in 8 attempts and succeeded in only one, completing the task in approximately 3 minutes. In contrast, OrionNav succeeded in all trials with low search times. This superior performance is attributed to the use of the open-world hierarchical scene graph and the LLM-based planner. The LLM Planner identifies the most probable room for the queried object from the scene graph and commands the system to search within that room, leading to successful object detection. In comparison, the Frontier-Search baseline lacks these capabilities, relying solely on frontier-based or random exploration to locate the object. Therefore, even in cases where dynamic changes in the environment prevent the object from being detected during the robot’s initial mapping, OrionNav equips the system with the capability to successfully locate the object in subsequent attempts.
@@ V-C2 Object-Map-Search
The key difference between this baseline and OrionNav is that the baseline uses the semantic object map without room labeling for object search. As a result, its success rate is approximately 56%, compared to OrionNav’s 100% success rate, highlighting the importance of the hierarchical scene graph in the proposed method. An example of this comparison is illustrated in Fig. 13, where the system is given a partial map of an environment that includes an office room, a break room, and a long L-shaped corridor. A monitor is dynamically added to the office room later, and the system is tasked with finding it. As shown in Fig. 13a, the baseline system hypothesizes that a monitor would be located on a table, so it navigates to a nearby table. However, this table was inside the break room and did not have the monitor. The baseline system then issues a global exploration command and eventually reaches the office room, where it detects the monitor. In contrast, OrionNav (Fig. 13b) leverages the hierarchical scene graph to infer that the monitor is likely in an office, navigating directly to it and detecting the monitor in a single step. Object-Map-Search, even in its successful cases, required more time to complete the task compared to OrionNav, as shown in Fig. 12.
@@ VI Conclusion
We proposed OrionNav, an autonomous system that integrates open-vocabulary hierarchical scene graphs with an LLM-based planner for real-time robot autonomy. The system builds a map of the environment using SLAM and fuses open-world semantics, room detection, and LLM-based labeling to generate a hierarchical scene graph. The LLM planner leverages this scene graph, along with user commands, to formulate high-level plans, which are executed by a lower-level safe controller. Extensive experiments in real-world environments demonstrated the system’s high success rates and practical applicability. Future research can explore mobile robots equipped with onboard manipulation capabilities, enabling tasks such as door opening and pick-and-place operations.

### v2

@@ V-C1 Frontier-Search
In our experiments, the Frontier-Search baseline demonstrated limited effectiveness, succeeding in only one out of nine trials while failing in the other eight attempts. The single successful trial required approximately three minutes to complete. In stark contrast, OrionNav achieved success in all trials with significantly reduced search times. This superior performance of our framework can be attributed to two key components: the open-world hierarchical scene graph and the LLM-based planner. OrionNav leverages the hierarchical scene graph to maintain a continuously updated understanding of the environment, including probable object locations based on semantic information. The LLM planner analyzes this scene graph to identify the room most likely to contain the queried object and efficiently directs the system to search within that specific area, leading to consistent and rapid object localization. Conversely, the Frontier-Search baseline lacks semantic understanding and relies solely on frontier-based or random exploration strategies to locate objects. This approach fails to adapt its exploration strategy based on the likelihood of object locations, resulting in inefficient searches and significantly lower success rates. Importantly, even when environmental changes prevent object localization during the robot’s initial mapping phase, OrionNav’s approach enables the system to successfully locate the object in subsequent attempts. This adaptability further underscores the robustness of OrionNav in complex, changing environments.
@@ V-C2 Object-Map-Search
The key distinction between the Object-Map-Search baseline and OrionNav lies in the former’s use of a semantic object map without room labeling for object search. This approach results in a success rate of about 56%, significantly lower than OrionNav’s 100% success rate, underscoring the importance of the hierarchical scene graph in the proposed method. Fig. 13 illustrates this comparison, presenting a scenario where both systems are provided with a partial map of an environment comprising an office room, a break room, and a long L-shaped corridor. The task involves locating a monitor that is temporarily removed from the office room for the initial mapping phase and then reintroduced to its original location. As shown in Fig. 13a, the baseline system assumes that a monitor would likely be located on a table and navigates to the nearest one. However, this table, situated in the break room, lacks a monitor. Consequently, it initiates a global exploration command to eventually reach the office room where it detects the monitor. In contrast, OrionNav (Fig. 13b) leverages its hierarchical scene graph to deduce that the monitor is likely in an office, navigating directly to it and detecting the monitor in a single step. As shown in Fig. 12, even in successful cases, the baseline consistently required more time to complete the task compared to OrionNav.
@@ VI Conclusion
We propose OrionNav, an autonomous system that integrates open-vocabulary hierarchical scene graphs with an LLM-based planner for real-time language-driven autonomous navigation. The system builds a map of the environment using SLAM and fuses open-world semantics, room detection, and LLM-based labeling to generate a hierarchical scene graph. The LLM planner leverages this scene graph, along with user commands, to formulate high-level plans, which are executed by a lower-level safe controller. The multi-level abstractions of perception and planning with continuously updated scene graphs and adaptive planner enables robust execution of diverse navigation tasks in large-scale complex, changing environments with the ability to recover from task failures. Extensive experiments in real-world environments demonstrated the system’s high success rates and practical applicability. Future research will explore mobile robots equipped with onboard manipulation and interaction capabilities, enabling tasks such as the opening of doors and pick-and-place.